# Peaceful Anarcho-Accelerationism: Decentralized Full Automation for a Society of Universal Care

### v3

@@ 5.1 Intersectional Liberation Through UDR
5.2 Maslow’s Hierarchy and the Liberation Stack
Abraham Maslow’s hierarchy of needs (Maslow, 1943), despite its well-known limitations as a rigid stage theory, captures a fundamental insight about human motivation: when basic survival needs are unmet, the higher capacities of creativity, intellectual exploration, moral reasoning, and self-actualization remain largely inaccessible. A hungry person does not philosophize; a person in physical danger does not compose music; a person consumed by the struggle for shelter does not pursue scientific inquiry. The hierarchy identifies five levels: physiological (food, water, shelter, warmth), safety (security, stability, health), belonging (love, friendship, community), esteem (recognition, contribution, competence), and self-actualization (creativity, purpose, spiritual growth).
The Liberation Stack (Section˜6) and UDR are designed, whether by explicit intention or structural convergence, to satisfy the first four levels universally, thereby liberating every human being to pursue the fifth. Figure˜6 illustrates the mapping.
@@ 5.2 Maslow’s Hierarchy and the Liberation Stack
Figure 6: Maslow’s hierarchy of needs mapped onto the Liberation Stack and UDR. The material layers of the stack (energy, manufacturing, food) satisfy physiological needs. UDR within ecological boundaries provides safety. Communication and governance commons enable belonging. Knowledge commons support esteem through contribution. Self-actualization, the pinnacle, is what remains when all other needs are met: creativity, spirituality, love, and the free development of each person’s unique potential.
The material layers of the Liberation Stack (energy, manufacturing, food) ensure that every person’s physiological needs are met: clean energy for warmth and cooking, automated manufacturing for shelter and clothing, agroecological production for nutritious food. UDR, constrained by ecological sustainability, provides safety: the security of knowing that one’s material needs will always be met, regardless of employment status, health, age, or any other contingency. Communication and governance commons provide belonging: the tools for meaningful participation in community life, for maintaining relationships across distance, for democratic self-governance. Knowledge commons provide esteem: the opportunity to contribute, to learn, to be recognized for one’s intellectual and creative work.
What remains, when the first four levels are universally satisfied, is self-actualization: the free development of each person’s unique potential. This is the society that anarcho-accelerationism envisions, not a society of idleness or consumption, but a society where every person is free to pursue the highest expressions of human existence: art, science, philosophy, spirituality, love, care, and the cultivation of community. This is real freedom.
5.3 The Transition Path: From Progressive UBI to UDR
A common objection to transformative visions is that they require an impossible leap from the current system to a radically different one. Anarcho-accelerationism rejects this false dichotomy. The transition from the current system to the care-centered society is progressive, and it leverages the very mechanisms of capitalist development to transcend capitalism itself.
The key insight is that automation is already happening, driven by the profit motive of capitalist firms. Every robot deployed in a factory, every AI system that replaces a white-collar function, every autonomous vehicle that eliminates a driving job, is a step toward full automation. Anarcho-accelerationism does not oppose this process; it redirects it. The fruits of capitalist technological development, the robots, the AI systems, the renewable energy infrastructure, the communication networks, are available for appropriation by the commons. This is not theft; these technologies were built on publicly funded research (Mazzucato, 2013), trained on collectively produced data, and developed using open-source tools created by volunteer communities.
The transition mechanism is a progressive continuum from Universal Basic Income (UBI) to Universal Desired Resources (UDR), illustrated in Figure˜7:
@@ 5.3 The Transition Path: From Progressive UBI to UDR
Figure 7: The progressive transition from the current system to full UDR. At each stage, workers displaced by automation are protected by progressively expanding mechanisms: first UBI (cash transfers within the market system), then UBI supplemented by growing commons infrastructure, and finally full UDR (direct access to the robotic commons). The transition leverages capitalist technological development while progressively building commons-based alternatives.
In the first stage, as automation displaces workers, governments introduce and progressively expand UBI, funded by taxation of automated production, carbon taxes, and wealth taxes. This is not anarcho-accelerationism’s endpoint but a transitional mechanism within the existing monetary system. UBI provides a floor of material security without challenging market dominance, yet it buys time and creates political space for the construction of commons infrastructure.
In the second stage, communities simultaneously build the Liberation Stack: community energy cooperatives, fablabs, agroecological food systems, mesh networks, knowledge commons, governance platforms. As commons infrastructure grows, an increasing share of material needs is met outside the market. The effective purchasing power of UBI increases because fewer goods need to be purchased, since energy comes from the community microgrid, food from the community farm, communication from the community mesh network. UBI begins to evolve from a monetary transfer into a guarantee of access to commons resources.
In the third stage, when the robotic commons can autonomously produce all material goods and services, money becomes unnecessary. UBI is replaced by UDR: every person requests what they need from the productive commons, subject only to ecological sustainability constraints. The transition is complete not through revolution but through the progressive construction of an alternative infrastructure that makes the old system obsolete.
This progressive path embodies what we term synthetic liberty, in contrast to the negative liberty of neoliberalism. Isaiah Berlin’s classic distinction (Berlin, 1969) between negative liberty (freedom from interference) and positive liberty (freedom to act, to realize one’s potential) illuminates the inadequacy of the neoliberal conception of freedom. Neoliberalism offers negative liberty: you are “free” to choose your employer, your health insurance, your housing, but only if you have the money to exercise that choice. A person who is formally free to choose but materially unable to afford any meaningful option is not free in any substantive sense (Sen, 1999). Van Parijs (1995) argues that real freedom requires not just the absence of coercion but the material conditions for genuine self-determination.
Synthetic liberty, as we define it, is the positive freedom that emerges when commons-based infrastructure provides the material conditions for genuine autonomy. It is “synthetic” in two senses: first, it is constructed (not naturally occurring; it must be built through deliberate institutional and technological design); second, it synthesizes the insights of multiple traditions (liberal autonomy, socialist equality, ecological sustainability, feminist care, communitarian belonging). Synthetic liberty is what remains when the material constraints that make negative liberty a hollow promise are removed by the robotic commons. It is the freedom to create, to care, to learn, to love, to worship, to build community, not because you have “chosen” these activities from a market menu, but because the material conditions of your existence no longer compel you to do otherwise.
6 The Liberation Stack: A Technical Architecture
Anarcho-accelerationism is not merely a philosophical position. It proposes a concrete layered architecture for sociotechnical systems that we term the Liberation Stack, illustrated in Figure˜8. Each layer is composed of existing, mature, open-source technologies.
@@ 6 The Liberation Stack: A Technical Architecture
Figure 8: The Liberation Stack: a layered architecture for sociotechnical sovereignty. The lower layers (0–2) provide material sovereignty over energy, manufacturing, and food. The upper layers (3–5) provide informational sovereignty over communication, knowledge, and governance. Each layer can be adopted independently, but the layers reinforce each other compositionally.
No community is truly autonomous if it depends on external entities for energy. Layer 0 provides energy sovereignty through community-owned solar photovoltaic arrays and wind turbines, battery storage systems, smart microgrid management via open-source controllers such as OpenEMS and Home Assistant, and peer-to-peer energy trading using transparent ledgers. Som Energia in Spain, a renewable energy cooperative with over 80,000 members, demonstrates that citizen-owned energy is commercially viable (Kunze and Becker, 2015). Brooklyn Microgrid in New York demonstrates peer-to-peer solar energy trading at the neighborhood scale (Mengelkamp et al., 2018).
Full automation of material production requires abundant, cheap, clean energy, and the accelerationist ecological thesis holds that the answer to the energy challenge is not degrowth but directed technological acceleration toward clean abundance (Jacobson et al., 2017). The convergence of multiple frontier energy technologies, all maturing simultaneously in the 2020s, makes this thesis empirically credible for the first time in history. We term this vision accelerationist ecologism: the conviction that ecological sustainability is best achieved not by reducing human activity but by powering it entirely with clean, renewable, and ultimately abundant energy sources that operate within planetary boundaries (Raworth, 2017).
Table˜6 summarizes the frontier energy technologies that underpin the energy layer of the Liberation Stack. These are not speculative proposals; they represent technologies with demonstrated prototypes, substantial investment, and clear pathways to commercial deployment.
Table 6: Frontier energy technologies for accelerationist ecologism: demonstrated innovations from 2020 onward that enable abundant clean energy for full automation.
Technology Key milestone (2020–2026) Projected impact Decentralization potential
Perovskite-silicon tandem solar 34.85% efficiency (LONGi, 2025); printable on flexible substrates (Green et al., 2023) Solar below \$0.01/kWh; panels on every surface Very high: modular, local manufacturing
Enhanced geothermal (EGS) Fervo Energy: first commercial 400 MW plant (2024–2026); 24/7 baseload (Tester et al., 2006) Geothermal anywhere on Earth; 5,000+ GW US potential High: community-scale plants
Nuclear fusion NIF ignition Q=4.13 (2025); CFS SPARC completion (2026); Helion/Microsoft PPA (Creely et al., 2020) Near-unlimited baseload energy by 2035–2040 Medium: regional plants, no carbon
Iron-air batteries Form Energy: 100+ hours storage at <<\$20/kWh; zero thermal runaway (Kwon et al., 2024) Solves multi-day renewable intermittency Very high: uses iron and air
Green hydrogen H2Pro E-TAC: 95%+ efficiency, \$1/kg target (Chatterjee et al., 2021) Replaces fossil fuels in industry, transport High: modular electrolyzers
Long-duration storage DOE \$1.76B for compressed air; gravity storage commercialized (U.S. Department of Energy, 2024) Grid-scale seasonal storage High: geological storage sites
Small modular reactors NuScale NRC-certified; Oklo fast reactor recycles nuclear waste Carbon-free 24/7 bridging technology Medium: factory-built, truck-delivered
@@
The convergence of these technologies implies a three-phase energy trajectory for the Liberation Stack. In the near term (2025–2030), perovskite tandem solar cells, enhanced geothermal systems, and iron-air batteries provide the technical foundation for community energy sovereignty at dramatically lower costs than today. In the medium term (2030–2040), first-generation commercial fusion plants, green hydrogen at grid parity, and long-duration storage systems enable regional energy abundance sufficient to power full industrial automation. In the long term (2040+), mature fusion technology provides near-unlimited clean baseload energy, while direct air capture powered by this abundance actively reverses historical carbon emissions, healing the atmosphere rather than merely stabilizing it.
Critically, the most decentralizable technologies, solar, enhanced geothermal, iron-air batteries, and green hydrogen, are precisely those that mature earliest. This means that community energy sovereignty does not depend on fusion or other centralized technologies; it can be achieved with technologies that are already commercially available or will be within five years. Fusion and SMRs serve as complementary technologies for energy-intensive applications (heavy industry, materials recycling, large-scale direct air capture) while community-scale renewables provide the distributed foundation. The result is an energy commons that is simultaneously abundant, clean, decentralized, and resilient: precisely what full automation requires, and precisely what the planet needs.
Distributed manufacturing eliminates dependence on global supply chains. Fablabs equipped with 3D printers, CNC routers, and laser cutters, combined with open-source hardware designs (OSHWA-certified) and circular materials systems such as Precious Plastic recycling, enable communities to produce locally what they need. The global network of 2,500+ fablabs (Fab Foundation, 2023) demonstrates that distributed manufacturing is already operational. WikiHouse provides open-source housing designs that can be assembled by non-specialists (Parvin, 2013). Modular, repairable design philosophies (FairPhone, Framework Laptop) extend product lifespans and reduce waste.
@@
Food sovereignty combines traditional ecological knowledge with computational tools: agroecological farming based on polycultures, composting, and biological pest control; precision agriculture via IoT sensors, drones, and open-source analytics; urban vertical farms and aquaponics; and the Open Food Network platform connecting local producers and consumers. Cuba’s agroecological revolution, born of necessity during the Special Period, demonstrated that small farmers can produce the majority of a nation’s food with minimal fossil fuel inputs (Altieri and Nicholls, 2012). FarmBot provides open-source precision agriculture for community gardens (FarmBot Inc., 2023).
Digital sovereignty requires control over communication infrastructure at every level: the physical layer through community mesh networks following the guifi.net model (40,000+ nodes) (Baig et al., 2015); the protocol layer through ActivityPub for federation and Matrix for real-time communication; the application layer through Mastodon, PeerTube, PixelFed, and Lemmy; and the privacy layer through end-to-end encryption by default (Signal protocol, Matrix E2EE). The Fediverse already hosts 10+ million users across thousands of independently operated servers (Zulli et al., 2020).
Knowledge commons ensure that information serves everyone: open-access publishing through arXiv, PLoS, and DOAJ; open educational resources through MIT OCW, Khan Academy, and P2PU; open data from government, scientific, and environmental sources; and open-source AI models trained on public data. arXiv itself, hosting 2.4+ million papers, demonstrates that open-access knowledge dissemination is not only possible but preferred by the scientific community. Wikipedia’s 60 million articles in 300+ languages represent the largest knowledge commons in human history.
Democratic self-governance is implementable as software: Pol.is for real-time consensus visualization in large groups (Small et al., 2021); Loomio for asynchronous, consent-based group decisions; Decidim for participatory budgeting, already adopted by Barcelona, Helsinki, and over 400 institutions worldwide (Barandiaran et al., 2024); LiquidFeedback for delegable, revocable, issue-specific voting; and ActivityPub-based inter-community coordination for federation across scales.
The Liberation Stack is designed for compositional sovereignty: each layer can be adopted independently, but the layers reinforce each other. A community with energy sovereignty has more bargaining power. A community with communication sovereignty cannot be silenced. A community with governance tools can make collective decisions about all other layers. Together, the layers form a complete alternative infrastructure that does not depend on either state provision or market availability. Table˜7 summarizes the architecture.
@@ 0 Energy Energetic sovereignty Solar cooperatives, OpenEMS, microgrids
Injustices Poverty Ecological crisis Techno-feudalism Precarity Intersectional oppression Understanding Reality Econometric evidence Scientific frontiers Historical analysis Commons research Intellectual Commitment Chomsky (Chomsky, 1967): “responsibility of intellectuals” Theory + praxis Social Levers Liberation Stack Open-source commons UBI $\to$ UDR Federated governance Peaceful transition MotivatesInformsActivatesTransforms: eliminates material basis of injustice

### v5

@@ 3 Bridge-Building and Intersectional Liberation
The structural insight is this: intersectional oppression is maintained through material scarcity administered through markets. When markets are superseded by the robotic commons under UDR, the material infrastructure of oppression is dismantled at its foundation. Cultural prejudices do not disappear automatically, but prejudice loses its institutional teeth: the capacity to deny resources, opportunities, and dignity on the basis of identity.
Abraham Maslow’s hierarchy of needs (Maslow, 1943) captures a fundamental insight about human motivation: when basic survival needs are unmet, the higher capacities of creativity, intellectual exploration, and self-actualization remain largely inaccessible. The Liberation Stack and UDR are designed to satisfy the first four levels universally, physiological needs through the energy, manufacturing, and food layers; safety through UDR within ecological boundaries; belonging through communication and governance commons; esteem through knowledge commons, thereby liberating every human being to pursue self-actualization. Figure˜5 illustrates the mapping.
@@
Figure 5: Maslow’s hierarchy of needs mapped onto the Liberation Stack and UDR. Self-actualization is what remains when all other needs are met: creativity, spirituality, love, and the free development of each person’s unique potential.
The transition from the current system to the care-centered society is progressive, not revolutionary, and it leverages the very mechanisms of capitalist development to transcend capitalism itself. Automation is already happening, driven by the profit motive. Every robot deployed, every AI system that replaces a function, is a step toward full automation. Anarcho-accelerationism does not oppose this process; it redirects it. The transition mechanism is a progressive continuum from Universal Basic Income to UDR, illustrated in Figure˜6. In the first stage, governments introduce and expand UBI, funded by taxation of automated production. In the second stage, communities simultaneously build the Liberation Stack: energy cooperatives, fablabs, agroecological food systems, mesh networks. As commons infrastructure grows, an increasing share of material needs is met outside the market. In the third stage, when the robotic commons can autonomously produce all material goods, money becomes unnecessary. UBI is replaced by UDR. The transition is complete not through revolution but through the progressive construction of an alternative infrastructure that makes the old system obsolete.
@@
Figure 6: The progressive transition from the current system to full UDR. At each stage, workers displaced by automation are protected by progressively expanding mechanisms.
This progressive path embodies what we term synthetic liberty, in contrast to the negative liberty of neoliberalism. Isaiah Berlin’s classic distinction (Berlin, 1969) between negative liberty (freedom from interference) and positive liberty (freedom to act) illuminates the inadequacy of the neoliberal conception of freedom. Van Parijs (1995) argues that real freedom requires the material conditions for genuine self-determination. Synthetic liberty is the positive freedom that emerges when commons-based infrastructure provides those conditions. It is synthetic in two senses: it is constructed through deliberate institutional and technological design, and it synthesizes the insights of multiple traditions.
4 The Liberation Stack, Accelerationist Ecologism, and the Pareto Frontier
Anarcho-accelerationism proposes a concrete layered architecture for sociotechnical systems that we term the Liberation Stack, illustrated in Figure˜7. Each layer is composed of existing, mature, open-source technologies.
@@ 4 The Liberation Stack, Accelerationist Ecologism, and the Pareto Frontier
Figure 7: The Liberation Stack: a layered architecture for sociotechnical sovereignty. Lower layers (0–2) provide material sovereignty. Upper layers (3–5) provide informational sovereignty. Layers reinforce each other compositionally.
No community is truly autonomous if it depends on external entities for energy. Layer 0 provides energy sovereignty through community-owned solar arrays and wind turbines, battery storage, smart microgrid management via open-source controllers such as OpenEMS, and peer-to-peer energy trading. Som Energia in Spain, a renewable energy cooperative with over 80,000 members, demonstrates that citizen-owned energy is commercially viable (Kunze and Becker, 2015). Brooklyn Microgrid demonstrates peer-to-peer solar energy trading at the neighborhood scale (Mengelkamp et al., 2018). Full automation requires abundant, cheap, clean energy, and the accelerationist ecological thesis holds that the answer is not degrowth but directed technological acceleration toward clean abundance (Jacobson et al., 2017). We term this vision accelerationist ecologism: the conviction that ecological sustainability is best achieved not by reducing human activity but by powering it entirely with clean, renewable, and ultimately abundant energy sources that operate within planetary boundaries (Raworth, 2017). Table˜7 summarizes the frontier energy technologies.
Table 7: Frontier energy technologies for accelerationist ecologism: demonstrated innovations enabling abundant clean energy for full automation.
Technology Key milestone Projected impact Decentralization potential
Perovskite-silicon tandem solar 34.85% efficiency (Green et al., 2023) Solar below \$0.01/kWh Very high: modular, local
Enhanced geothermal Fervo Energy: 400 MW plant (Tester et al., 2006) Geothermal anywhere on Earth High: community-scale
Nuclear fusion NIF ignition Q=4.13; SPARC 2026 (Creely et al., 2020) Near-unlimited baseload by 2035–2040 Medium: regional plants
Iron-air batteries Form Energy: 100+ hours, <<\$20/kWh (Kwon et al., 2024) Solves multi-day intermittency Very high: iron and air
Green hydrogen H2Pro: 95%+ efficiency (Chatterjee et al., 2021) Replaces fossil fuels in industry High: modular electrolyzers
Long-duration storage DOE \$1.76B investment (U.S. Department of Energy, 2024) Grid-scale seasonal storage High: geological sites
@@
Distributed manufacturing eliminates dependence on global supply chains. The global network of 2,500+ fablabs (Fab Foundation, 2023) demonstrates that distributed manufacturing is operational. WikiHouse provides open-source housing designs (Parvin, 2013). Food sovereignty combines traditional ecological knowledge with computational tools: agroecological farming, precision agriculture via IoT sensors and drones, and the Open Food Network connecting local producers and consumers. Cuba’s agroecological revolution demonstrated that small farmers can produce the majority of a nation’s food with minimal fossil fuel inputs (Altieri and Nicholls, 2012).
Digital sovereignty requires control over communication infrastructure at every level: the physical layer through community mesh networks following the guifi.net model with 40,000+ nodes (Baig et al., 2015); the protocol layer through ActivityPub for federation and Matrix for real-time communication; the application layer through Mastodon, PeerTube, PixelFed, and Lemmy; and the privacy layer through end-to-end encryption. The Fediverse hosts over 10 million users across thousands of independently operated servers, governed by their communities without centralized control (Zulli et al., 2020). Its governance model maps directly onto Ostrom’s design principles: clear boundaries (each server defines its rules), proportional equivalence (operators bear costs proportional to use), collective choice (community governance of moderation), monitoring (transparent moderation logs), graduated sanctions (muting, suspension, defederation), conflict resolution (inter-server negotiation), minimal recognition of rights (ActivityPub as W3C standard), and nested enterprises (individual servers within the broader network). The contrast with centralized platforms is instructive: when Elon Musk acquired Twitter in 2022, a single individual’s decisions could transform the governance of a platform used by hundreds of millions, an event that would be structurally impossible in a federated architecture. Knowledge commons ensure that information serves everyone: arXiv hosts 2.4+ million papers, Wikipedia’s 60 million articles in 300+ languages represent the largest knowledge commons in human history. Democratic self-governance is implementable as software: Pol.is for consensus visualization (Small et al., 2021), Decidim for participatory budgeting (Barandiaran et al., 2024), LiquidFeedback for delegable voting.
@@ 0 Energy Energetic sovereignty Solar cooperatives, OpenEMS, microgrids
The accelerationist ecological thesis is supported by rigorous econometric evidence. The Environmental Kuznets Curve hypothesis (Grossman and Krueger, 1995; Panayotou, 1993) proposes an inverted-U relationship between income and environmental degradation. The most comprehensive systematic review, screening 11,500+ papers, confirms that absolute decoupling of GDP from emissions has been achieved in several developed economies (Haberl et al., 2020). Le Quéré et al. (2019) found that renewable energy deployment is the primary driver of sustained emissions reductions in 18 developed economies. Way et al. (2022) demonstrate that a fast transition to renewables would save approximately \$12 trillion compared to maintaining fossil fuels. Acemoglu et al. (2012) provide the theoretical foundation for directed technical change: policy interventions can permanently redirect innovation toward clean technologies. Combined with dematerialization (McAfee, 2019), the evidence strongly supports the accelerationist thesis.